The tires of a car make 83 revolutions as the car reduces its speed uniformly from 90.0 km/h to 63.0 km/h. The tires have a diameter of 0.82 m. Part A What was the angular acceleration of the tires? Express your answer using two significant figures. α = rad/s2 Part B If the car continues to decelerate at this rate, how much more time is required for it to stop? Express your answer to two significant figures and include the appropriate units. t = nothing nothing Request Answer Part C If the car continues to decelerate at this rate, how far does it go? Find the total distance. Express your answer to three significant figures and include the appropriate units.
Answer #1

a)
Initial velocity, u = 90 km/h
= 90 * 5/18 m/s
= 25 m/s
Initial angular velocity, i = u/R
Where R is the radius of the wheel.
= 25/0.82 = 30.49 rad/s

Final velocity, v = 63 m/s
= 63 * 5/18 m/s
= 17.5 m/s
FInal angular velocity, f = v/R
= 17.5 / 0.82 = 21.34 rad/s

Angular distance travelled, = 83 rev
= 83 * 2 rad
= 521.5 rad

Using the formula, f2 - i2 = 2
= (f2 - i2) / 2
= (21.342 - 30.492) / (2 * 521.5)
= - 0.45 rad/s2

b)
Using the formula, f = i + t,
Where is the angular acceleration and t is the time taken.
t = (f - i) /
= (21.34 - 30.49) / 0.45
= 20 s

c)
Initial angular speed = f = 21.34 rad/s (from part a)
Final angular speed, f' = 0
Angular acceleration, = - 0.45 rad/s2
Consider as the ' as the angular displacement

Using the formula, f'2 - f2 = 2'
' = (f'2 - f2) / 2
= (0 - 21.342) / (2 * 0.45)
= 501 rad

Total distance = + '
= 521.5 + 501
= 1022.6 rad = 163 revolutions
= 1020 rad = 163 rev